OpenStudy (anonymous):

2x=5+4x solve by completing the square.

OpenStudy (anonymous):

there is no need of completing square it is linear variable can be solved easily put all x terms on one side and all constants on other side 4x-2x=-5 2x=5 x=5/2

OpenStudy (anonymous):

4x-2x=-5 2x=-5 x=-5/2

OpenStudy (anonymous):

ERR SORRY I TYPED IT WRONG. 2x=5+4/x solve by completing the square.

OpenStudy (anonymous):

yeahh i agree with sami!

OpenStudy (anonymous):

guys?

OpenStudy (anonymous):

So 2x^2 = 5x +4 then...

OpenStudy (anonymous):

(x not zero obviously)

OpenStudy (anonymous):

uhhokay then... 2x^2-5x-4=0 right?

OpenStudy (anonymous):

then 2x^2-5x+(5/2)^2=4+(5/2)^2 ?

OpenStudy (anonymous):

Might as well leave the 4 on the rhs 2x^2 -5x = 4 And divide by 2 so we get a single x^2 x^2 -5/2 x = 2

OpenStudy (anonymous):

So you take half the coefficient of x.....

OpenStudy (anonymous):

o.o i have never done that before o.o

OpenStudy (anonymous):

so now what do we do?

OpenStudy (anonymous):

(x-5/4)^2 - (5/4)^2 = 2 etc...

OpenStudy (anonymous):

Just the usual from here....

OpenStudy (anonymous):

wait, i'm so confused. x^2-5/2x+(-5/4x)^2=2

OpenStudy (anonymous):

It's (x-5/4)^2 - (5/4)^2 = 2 (not what you have put)

OpenStudy (anonymous):

how do you get that though?

OpenStudy (anonymous):

By putting (x-5/4)^2 I have added (5/4)^2 to the left side so I have to take it away again to get back to where I started.